If x 2 +3x+3=0 and ax 2 +bx+1=0 ,(a,b belong to rational) have a common root, then value of (3a+b) is equal to : (A) 1/3 (B) 1 (C) 2 (D) 4

If x2+3x+3=0 and ax2+bx+1=0 ,(a,b belong to rational) have a common root, then value of (3a+b) is equal to :
(A) 1/3
(B) 1
(C) 2
(D) 4

Grade:11

1 Answers

Arun
5 years ago
1st equation has D 0
hence it has both roots imaginary 
Hence both equations have oth roots common
so
1/a = 3/b =3/1
a = 1/3
b = 1
hence
a/b = 1/3